1. A 25 Ω resistor and a 35 Ω  resistor are connected in series and placed across a 120 V source.  (a)  Determine the effective resistance of the circuit.  (b)  What is the current in the circuit?  (c)  What is the voltage drops across the 25 Ω resistor?  (d)  What is the power used in the 35 Ω resistor?
1
Expert's answer
2021-03-18T14:02:12-0400

Given:

A circuit with:

  • 2 resistors of 35 "\Omega" and 25 "\Omega"
  • resistors connected in series combintion
  • a source (battery) of 120 V

a) Effective Resistance of the circuit.

Let, R1=35"\Omega" and R2 = 25"\Omega"

As we know that the resistors are in series combination,

then current flow through both the resistors will be same.

"\therefore" equivalent resistance of the circuit will be given by

"\boxed{R_eq=R_1+R_2}"

"{R_eq= 35+ 25}"

"\boxed{R_eq=60}"



b) Current in circuit


According to OHM'S LAW Current in circuit is given by:

"\boxed {I={V\over R_eq}}"

here; V=120

Req=60


"\boxed{\therefore I={120\over60}=2 A}"


c) Voltage drop across 25 "\Omega" resistance

According to Ohm's law

in a series combination voltage drop across any resistor is given by

"\boxed{V_1=IR_1}"

here;

I=2A

R=25"\Omega"

"\therefore" voltage across 25 "\Omega" is

"\boxed{V_1=2 *25= 50 volts}"


d) power used in 35"\Omega" resistor

In a series combination;

Power used by load or resistor is given by :

"\boxed{P_1=I^2R_1}"

here,

I=2 A

R= 35"\Omega"

"\boxed{\therefore P=(2)^2 *35= 140 W}"